an airplane can fly 390 miles with the wind in the same amount of time it would take to fly 210 miles against the wind. If the plane fly 100 miles per hour is still air, what is the speed of the wind?
Found 2 solutions by mananth, richwmiller:
Answer by mananth(16946)

Let the speed of the wind be x mph
speed in still air = 100mph
speed against = 100-x
speed with wind = 100+x
time taken is same
390/100+x= 210/100-x
390(100-x)=210(100+x)
39000-390x=21000+210x
39000-21000=210x+390x
18000 = 600 x
x=18000/600
30 mph which is the speed of the wind

Answer by richwmiller(17219)

(100+w)*t=390
(100-w)*t=210
t=390/(100+w)
t=210/(100-w)
390/(100+w)=210/(100-w)
w=30
130*t=390
70*t=210
t=3